OpenStudy (anonymous):

Your grade consists of three 100 point exams and a final that is worth 150 points. You score a 71, 74, and 92 on the three exams. Your final grade in the class is 82%. How many points did you earn on the final?

OpenStudy (ash2326):

@justine97 Final grade is the percentage of score obtained . Suppose I score x in the final which is worth 150 points. Let's find my total score Can you find my total score ?

OpenStudy (anonymous):

So (71+74+92+x)/4 = 82?

OpenStudy (ash2326):

It will be divided by (100+100+100+150) and multiplied by 100 to find %

OpenStudy (ash2326):

do you follow?

OpenStudy (anonymous):

I'm not really sure what you mean by 100+100+100+150?

OpenStudy (ash2326):

This is the max score, three tests are for 100 max marks and one test for 150. So total 100+100+100+150

OpenStudy (anonymous):

Okay so how do you find out what you got on the final?

OpenStudy (ash2326):

How much total did you obtain?

OpenStudy (anonymous):

I got 136.5 points

OpenStudy (ash2326):

We do need to find the average. Just find the sum. Let's say that S \[S=71+74+92+x\]

OpenStudy (ash2326):

@justine97 Can you find the sum?

OpenStudy (anonymous):

How would I find that?

OpenStudy (ash2326):

S=71+74+92+x S=145+92+x S=237+x Do you follow?

OpenStudy (anonymous):

Yes.

OpenStudy (ash2326):

okay, let's find the total of max marks of all 4 tests. Can you find that?

OpenStudy (anonymous):

I got 450 pts. What do we do next?

OpenStudy (ash2326):

now \[\frac{237+x}{450}\times 100=82\] Can you find x from here?

OpenStudy (anonymous):

I got 132. Is that the right answer?

OpenStudy (ash2326):

correct, do you understand?

OpenStudy (anonymous):

Yes. Thank you!! :')